OpenStudy (anonymous):

Why is my answer to this ap calc question completely completely off? :( help please! Thank you!!

OpenStudy (anonymous):

what is the question? what is your answer?

OpenStudy (anonymous):

oops Ill attached the photo

OpenStudy (anonymous):

Do you see the question now?

OpenStudy (anonymous):

There's an error in your fourth line: \[200=-\frac{6.89}{0.053}e^0+C~~\implies~~200=-\frac{6.89}{0.053}+C~~\implies~~C=330\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

But then I get 224 for my answer and its not right?

OpenStudy (anonymous):

Oh I didn't notice the other mistake you made earlier. When you integrated to find \(T(t)\), you have two negatives that should cancel out: \[-6.89\int e^{-0.053t}\,dt=\frac{-6.89}{-0.053}e^{-0.053t}+C=\color{red}{130}e^{-0.053t}+C\]

OpenStudy (anonymous):

ohh right thank you! and so C comes out to be 70, and is that why the question tells you to room temperature is 70 degrees on average? Because I thought that was just a trick in the question as we are dealing with tea temp not room temp..but why would that just be the constant?

OpenStudy (anonymous):

Well as the tea cools indefinitely, it will reach the ambient temperature. Notice that as \(t\to\infty\), the exponential term approaches \(0\), so you're left with \(70^\circ\), the room temperature.